OpenStudy (anonymous):

More Trigonometric Proofs \[2\sec^2=\frac{ 1-\sin t }{ \cos^2 t} + \frac{ 1 }{ 1-sint }\]

OpenStudy (anonymous):

Do I multiply by the conjugate?

OpenStudy (misty1212):

HI (again)

OpenStudy (misty1212):

i would add

OpenStudy (anonymous):

Hi again xD

OpenStudy (anonymous):

So find common denominator?

OpenStudy (misty1212):

not really "find" one, just multiply the denominators together

OpenStudy (anonymous):

Yeah same thing

OpenStudy (anonymous):

oh so one of the fractions reduces to 1

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

Oh wait, oops

OpenStudy (misty1212):

here is what i do it is too confusing to do the algebra with sines and cosines , so i would replace cosine by \(a\) and sine by \(b\) and add up \[\frac{1-b}{a^2}+\frac{1}{1-b}\]

OpenStudy (anonymous):

so now I have \[\frac{ \cos^2(1-\sin t) }{ \cos^2t(1-sint) } + \frac{ \cos t }{ \cos^2t(1-sint)}\]

OpenStudy (misty1212):

yeah but it is a lot easier to write \[\frac{(1-b)(1-b)+a^2}{a^2(1-b)}\]

OpenStudy (anonymous):

Sorry left side on numerator is cos^2(1-sin^2t)

OpenStudy (misty1212):

yeah i know but it is annoying to write \[\frac{ \cos^2(1-\sin t) }{ \cos^2t(1-sint) } + \frac{ \cos^2 t }{ \cos^2t(1-sint)}\]

OpenStudy (anonymous):

Yeah okay

OpenStudy (anonymous):

So then combine

OpenStudy (misty1212):

\[\frac{(1-b)(1-b)+a^2}{a^2(1-b)}\] is a lot easier now lets work from there

OpenStudy (misty1212):

multiply out and get \[\frac{1-2b+b^2+a^2}{a^2(1-b)}\]

OpenStudy (misty1212):

now comes one trig step, namely \(a^2+b^2=1\) because \(\cos^2(x)+\sin^2(x)=1\)

OpenStudy (misty1212):

you get \[\frac{2-2b}{a^2(1-b)}\]

OpenStudy (anonymous):

Oh okay, so now it becomes 2-2b

OpenStudy (anonymous):

Okay

OpenStudy (misty1212):

factor out the 2, cancel the \(1-b\) top and bottom and voila

OpenStudy (anonymous):

ah okay! Thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (wade123):

@satellite73 hii can you please come continue with me after youre done here? its lagging so i dont kow if youre getting my notifications